OpenStudy (anonymous):

Decide whether the equation is a trigonometric identity. Explain your reasoning. 1. cos^2 x (1+tan^2 x) =1

OpenStudy (anonymous):

i got yes

OpenStudy (cwrw238):

sec^2 x = 1 + tan^2 x so substituting: = cos^2 x * sec^2 x but sec^2 x = 1/cos^ 2 so - can you see the way to go?

OpenStudy (anonymous):